OpenStudy (anonymous):

HELP !! Giving Out Medals ;) NEED ANSWER NOW. Write an equation for the line that is parallel to the given line and that passes through the given point. y=3/4x-9; (-8,-18) A. y=4/3x-12 B. y=3/4x-12 C. y=-4/3x+12 D. y=3/4x+11/2

OpenStudy (anonymous):

parallel lines have same slope so m = 3/4 using point slope form y -(-18) = 3/4(x - (-8)) y +18 = 3/4x +6 y = 3/4x +6 -18 y = 3/4x -12